OpenStudy (anonymous):

(solve, medal given) f(-3)= 2(-3)^3 + 4(-3)^2 + (-3)-1

OpenStudy (anonymous):

what are you trying to solve

OpenStudy (anonymous):

Ur Mom

OpenStudy (anonymous):

= 2(-27) + 4(9) + 3 = -54 + 36 + 3 = -15

OpenStudy (anonymous):

if you're trying to simplify then it would be -15